OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution.

OpenStudy (anonymous):

\[3-n \over 2 \] = \[3 \over 4\]

OpenStudy (anonymous):

In other words, 3-n/2 = 3/4

OpenStudy (turingtest):

1)multiply both sides by four 2)collect like terms 3)isolate n

OpenStudy (ash2326):

(3-n)/2=3/4 multipying 4 to both sides of equation 4(3-n)/2=4*3/4 2(3-n)=3 expand the bracket 6-2n=3 subtract 6 from both sides 6-2n-6=3-6 -2n=-3 so n=1.5

OpenStudy (anonymous):

are you good to go? Or do you still need help?